# A hierarchical approach for multicontinuum homogenization in high contrast media

### v2

@@ 3.4 Computational savings
Now, we compare the computational complexity of the original multicontinuum homogenization method and our hierarchical algorithm. Again, let $h$ denote the FE mesh size for the constraint problems (8), $H$ denote the initial coarse mesh size, and $\eta$ be the coarsening factor. In the hierarchical approach, for the $n$-th level, the degrees of freedom of the local problems at one macropoint are $\mathcal{O}\left((\eta^{n-1}h)^{-d}\right)$, while the number of $n$-th macropoints is $\mathcal{O}\left((\eta^{L-n}H)^{-d}\right)$. The computational complexity for one level is $\mathcal{O}\left((hH)^{-d}\eta^{(1-L)d}\right)$. Since there are $L$ levels, the total computational cost is $\mathcal{O}\left(L(hH)^{-d}\eta^{(1-L)d}\right)$. In contrast, for the original multicontinuum homogenization method, the degrees of freedom for local problems at one macropoint are $\mathcal{O}(h^{-d})$, while the number of macropoints is $\mathcal{O}(H^{-d})$. Thus, the total computational cost is $\mathcal{O}\left((hH)^{-d}\right)$. Note that $\mathcal{O}(L\eta^{(1-L)d})<\mathcal{O}(1)$, which highlights the significant reduction in computational cost achieved by our hierarchical approach.
@@ 4 Numerical examples
In this section, we present four numerical examples to demonstrate the efficiency of our method. Our codes are available on GitHub [^1] https://github.com/xieweidc/hierarchical_mh_elliptic.git. The permeability field is defined as $\kappa(x)=g_{1}(x)g_{2}(x)$. Here, $g_{1}(x)$ is a slowly varying function given by $g_{1}(x)=2+\sin(\pi x_{1})\sin(\pi x_{2})$, and $g_{2}(x)$ is a high-contrast function in two different regions, detailed as follows:
[…]
where $\Omega_{1}$ and $\Omega_{2}$ represent the low and high conductivity regions, respectively. The geometry configuration will be shown in the specific examples. For clarity, we will specify the parameter $\epsilon$ associated with each example individually. The source term is denoted as
$f(x)=\begin{cases}\begin{aligned} \frac{\epsilon}{10}e^{-40|(x_{1}-0.5)^{2}+(x_{2}-0.5)^{2}|},\quad&x\in\Omega_{1},\\ e^{-40|(x_{1}-0.5)^{2}+(x_{2}-0.5)^{2}|},\quad&x\in\Omega_{2}.\end{aligned}\end{cases}$
The computational domain $\Omega$ is defined as a unit square. The fine-grid solution is obtained by using a rectangular partition with a mesh size of $h=1/960$. We take the whole coarse element as an RVE, namely, $R_{p}=K_{p}$. Additionally, we set the oversampling layers $l$ to be $\lceil-2\log(H)\rceil$ for all the tests. The relative $L^{2}$ error for each continuum is defined as follows:
$e_{2}^{(i)}=\sqrt{\cfrac{\sum_{p}\left|\frac{1}{|K_{p}|}\int_{K_{p}}U_{i}-\frac{1}{|{K_{p}}\cap\Omega_{i}|}\int_{{K_{p}}\cap\Omega_{i}}u\right|^{2}}{\sum_{p}\left|\frac{1}{|{K_{p}}\cap\Omega_{i}|}\int_{{K_{p}}\cap\Omega_{i}}u\right|^{2}}}.$
In our paper, we compute three different types of errors:
  • Type 1: The error of multicontinuum homogenization using the fine-grid solution as the reference solution. This error primarily investigates the performance of the original multicontinuum homogenization method.
  • Type 2: The error of hierarchical multicontinuum homogenization using the fine-grid solution as the reference solution. This error evaluates the performance of the hierarchical multicontinuum homogenization method.
  • Type 3: The error of hierarchical multicontinuum homogenization using the multicontinuum homogenization solution as the reference solution. This error studies the difference between the hierarchical algorithm and the fully high-resolution computation.
@@ 4.1 Example 1
In this example, we consider the high-contrast function $g_{2}$ as a layered medium with $\epsilon=1/48$. The logarithm of the function $\log(g_{2})$ and the reference solution are depicted in Figure 5. In Table 1, we observe that three types of errors consistently decrease as the coarse mesh size is refined. Notably, the difference between the multicontinuum homogenization method and the hierarchical multicontinuum homogenization method remains negligible. For instance, when $H=1/24$, both methods yield errors on the order of $10^{-2}$, with a difference between the two approaches as small as $10^{-6}$. In Figure 6, we depict the average solution of the reference solution, the multicontinuum homogenization method, and the hierarchical scheme. Visually, no discernible differences are evident, confirming the efficacy of the proposed methods.
Figure 5: High-contrast part $\log(g_{2})$ (left) and reference solution (right) in example 1.
@@
$\frac{1}{12}$ $\lceil-2\log(H)\rceil=5$ 8.17e-02 7.60e-02 8.17e-02 7.60e-02 1.36e-06 5.45e-07
$\frac{1}{24}$ $\lceil-2\log(H)\rceil=7$ 1.84e-02 1.49e-02 1.84e-02 1.49e-02 1.30e-07 7.64e-09
$\frac{1}{48}$ $\lceil-2\log(H)\rceil=8$ 2.76e-03 2.31e-03 2.76e-03 2.31e-03 4.27e-08 3.56e-08
@@
Figure 6: Average solution in example 1. Top-Left: Average reference solution in $\Omega_{1}$. Top-Right: Average reference solution in $\Omega_{2}$. Middle-Left: Average multicontinuum homogenization solution in $\Omega_{1}$. Middle-Right: Average multicontinuum homogenization solution in $\Omega_{2}$. Bottom-Left: Average hierarchical multicontinuum homogenization solution in $\Omega_{1}$. Bottom-Right: Average hierarchical multicontinuum homogenization solution in $\Omega_{2}$.
@@ 4.2 Example 2
In this example, we examine the high-contrast function $g_{2}$ with a crossed geometry, characterized by $\epsilon=1/48$. We depict the logarithm of the function $g_{2}$ and the reference solution in Figure 7. The errors obtained by refining the coarse mesh size are detailed in Table 2. We observe that our hierarchical algorithm has the same performance as the original multicontinuum homogenization method. The error decreases as the coarse mesh size decreases. Furthermore, the difference between these two approaches is negligible compared to the error stemming from the homogenization process. For example, the error due to homogenization can be as large as $10^{-3}$, while the difference between the two methods can be as small as $10^{-8}$ when $H=1/48$. To provide additional validation, we select representative average solutions, depicted in Figure 8, which further demonstrate the efficiency of our method.
Figure 7: High-contrast part $\log(g_{2})$ (left) and reference solution (right) in example 2.
@@
$\frac{1}{12}$ $\lceil-2\log(H)\rceil=5$ 4.54e-02 4.53e-02 4.54e-02 4.53e-02 2.75e-06 2.52e-06
$\frac{1}{24}$ $\lceil-2\log(H)\rceil=7$ 7.07e-03 6.22e-03 7.07e-03 6.22e-03 2.37e-07 1.43e-07
$\frac{1}{48}$ $\lceil-2\log(H)\rceil=8$ 1.76e-03 1.40e-03 1.76e-03 1.40e-03 6.78e-08 6.05e-08
@@
Figure 8: Average solution in example 2. Top-Left: Average reference solution in $\Omega_{1}$. Top-Right: Average reference solution in $\Omega_{2}$. Middle-Left: Average multicontinuum homogenization solution in $\Omega_{1}$. Middle-Right: Average multicontinuum homogenization solution in $\Omega_{2}$. Bottom-Left: Average hierarchical multicontinuum homogenization solution in $\Omega_{1}$. Bottom-Right: Average hierarchical multicontinuum homogenization solution in $\Omega_{2}$.
@@ 4.3 Example 3
In the third example, we consider a non-periodic high-contrast function $g_{2}$, with a contrast parameter of $\epsilon=1/30$. For clarity, we plot the logarithm of $g_{2}$ and the fine-grid solution in Figure 9. Accurately representing the different continua requires a very fine grid, and the non-periodic case further increases this demand. Due to the limitations of our computational resources, we test only two different coarse mesh sizes for the non-periodic structure. The results are detailed in Table 3. We observe that the homogenization error decreases as the coarse mesh size decreases. However, the hierarchical part failed to maintain the same property, but it can provide an acceptable accuracy, achieving a $10\%$ error for a given mesh size. In Figure 10, we plot the three different average solutions with $H=1/24$. From the figure, we observe that this is consistent with our error table.
Figure 9: High-contrast part $\log(g_{2})$ (left) and reference solution (right) in example 3.
@@
$\frac{1}{12}$ $\lceil-2\log(H)\rceil=5$ 4.74e-02 4.70e-02 1.45e-01 1.03e-01 1.03e-01 6.11e-02
$\frac{1}{24}$ $\lceil-2\log(H)\rceil=7$ 1.33e-02 1.07e-02 1.20e-01 8.06e-02 1.22e-01 8.37e-02
@@
Figure 10: Average solution in example 3. Top-Left: Average reference solution in $\Omega_{1}$. Top-Right: Average reference solution in $\Omega_{2}$. Middle-Left: Average multicontinuum homogenization solution in $\Omega_{1}$. Middle-Right: Average multicontinuum homogenization solution in $\Omega_{2}$. Bottom-Left: Average hierarchical multicontinuum homogenization solution in $\Omega_{1}$. Bottom-Right: Average hierarchical multicontinuum homogenization solution in $\Omega_{2}$.
@@ 4.4 Example 4
In the last example, we consider another non-periodic configuration of $g_{2}$. The contrast parameter $\epsilon$ is $1/12$. The distribution of the two continua and the reference solution is depicted in Figure 11. For the same reason as in Example 3, we refine the coarse mesh only once. In Table 4, we provide details of the numerical errors. We observe that the error in the homogenization part decreases as we refine the coarse mesh. The error in the hierarchical part does not change but is kept within an acceptable tolerance, around $10\%$. In Figure 12, we show the average solutions of the fine-grid, multicontinuum homogenization, and hierarchical multicontinuum homogenization approaches. These figures are also consistent with our observations from the error table.
Figure 11: High-contrast part $\log(g_{2})$ (left) and reference solution (right) in example 4.
@@
$\frac{1}{12}$ $\lceil-2\log(H)\rceil=5$ 6.26e-02 6.41e-02 1.46e-01 1.33e-01 9.06e-02 7.54e-02
$\frac{1}{24}$ $\lceil-2\log(H)\rceil=7$ 1.95e-02 2.26e-02 1.15e-01 1.02e-01 1.07e-01 9.44e-02
@@
Figure 12: Average solution in example 4. Top-Left: Average reference solution in $\Omega_{1}$. Top-Right: Average reference solution in $\Omega_{2}$. Middle-Left: Average multicontinuum homogenization solution in $\Omega_{1}$. Middle-Right: Average multicontinuum homogenization solution in $\Omega_{2}$. Bottom-Left: Average hierarchical multicontinuum homogenization solution in $\Omega_{1}$. Bottom-Right: Average hierarchical multicontinuum homogenization solution in $\Omega_{2}$.
@@ Acknowledgement
Wei Xie acknowledges support from the China Scholarship Council (CSC, Project ID: 202308430231) for funding during the visit to Nanyang Technological University (NTU). Sincere thanks to NTU for providing an outstanding working and learning environment that greatly advanced the research and professional development. Viet Ha Hoang is supported by the Tier 2 grant T2EP20123-0047 awarded by the Singapore Ministry of Education. Yin Yang is supported by the National Natural Science Foundation of China Project (No. 12071402, No. 12261131501), the Project of Scientific Research Fund of the Hunan Provincial Science and Technology Department (No. 2024JJ1008), and Program for Science and Technology Innovative Research Team in Higher Educational Institutions of Hunan Province of China. Yunqing Huang is supported by National Natural Science Foundation of China Key Program (12431014).

### v6

@@ 3.4 Computational savings
Now, we compare the computational complexity of the original multicontinuum homogenization method and our hierarchical algorithm. Again, let $h$ denote the finest FE mesh size for the constraint problems (8), $H$ denote the initial coarse mesh size, and $\eta$ be the coarsening factor. In the hierarchical approach, for the $n$-th level, the degrees of freedom of the local problems at one macropoint are $\mathcal{O}\left((\eta^{n-1}h)^{-d}\right)$, while the number of $n$-th macropoints is $\mathcal{O}\left((\eta^{L-n}H)^{-d}\right)$. The computational complexity for one level is $\mathcal{O}\left((hH)^{-d}\eta^{(1-L)d}\right)$. Since there are $L$ levels, the total computational cost is $\mathcal{O}\left(L(hH)^{-d}\eta^{(1-L)d}\right)$. In contrast, for the original multicontinuum homogenization method, the degrees of freedom for local problems at one macropoint are $\mathcal{O}(h^{-d})$, while the number of macropoints is $\mathcal{O}(H^{-d})$. Thus, the total computational cost is $\mathcal{O}\left((hH)^{-d}\right)$. Note that $\mathcal{O}(L\eta^{(1-L)d})<\mathcal{O}(1)$, which highlights the significant reduction in computational cost achieved by our hierarchical approach.
@@ 4 Numerical examples
In this section, we present five numerical examples to demonstrate the efficiency of our method, including both 2D and 3D cases. The permeability field is defined as $\kappa(x)=k(x)g(x)$, where $g(x)$ represents a slowly varying function over the global domain: $g(x)=2+\sin(\pi x_{1})\sin(\pi x_{2})$ for 2D cases, and $g(x)=e^{(x_{1}+x_{2}+x_{3})/2}$ for the 3D case. The function $k(x)$ describes a high-contrast distribution across two continua, with $k(x)=\eta$ in the low-conductivity region $\Omega_{1}$ and $k(x)=1$ in the high-conductivity region $\Omega_{2}$. We set $\eta=10^{-4}$ for 2D examples and $\eta=10^{-2}$ for the 3D case. The specific configurations of $k(x)$ are provided in the corresponding examples.
The source term is defined as
$f(x)=\begin{cases}\eta\,e^{-40\left((x_{1}-0.5)^{2}+(x_{2}-0.5)^{2}\right)},&x\in\Omega_{1},\\ e^{-40\left((x_{1}-0.5)^{2}+(x_{2}-0.5)^{2}\right)},&x\in\Omega_{2}.\end{cases}$
The computational domain is $\Omega=[0,1]^{d}$. For the 2D examples, the fine-grid solution is obtained using a rectangular mesh of size $h=1/960$, while for the 3D example, a hexahedral mesh with $h=1/120$ is employed. Each coarse element is treated as a representative volume element (RVE), i.e., $R_{p}=K_{p}$, and the number of oversampling layers is set to $l=\lceil-2\log(H)\rceil$ unless otherwise specified. All numerical experiments were carried out on an HPC cluster node equipped with dual AMD EPYC 9654 CPUs (192 physical cores, 384 threads) and 512 GB of memory.
To quantify the performance of the upscaling techniques, we define the average fine-grid solution over each coarse block as
$U_{h}^{(i)}=\frac{1}{|K_{p}\cap\Omega_{i}|}\int_{K_{p}\cap\Omega_{i}}u,$ (14)
and the corresponding multiscale solutions as
$U_{\textnormal{MH}}^{(i)}=\frac{1}{|K_{p}|}\int_{K_{p}}U_{i,\textnormal{MH}},\quad U_{\textnormal{HMH}}^{(i)}=\frac{1}{|K_{p}|}\int_{K_{p}}U_{i,\textnormal{HMH}},$ (15)
where “MH” and “HMH” denote the multicontinuum homogenization and hierarchical multicontinuum homogenization methods, respectively.
To compare two different average solutions for each continuum, we define
$e_{2}^{(i)}(U,V)=\sqrt{\frac{|U-V|^{2}}{|U|^{2}}}.$
In this work, we consider three types of errors:
  • Type 1: $e_{2}^{(i)}(U_{h},U_{\textnormal{MH}})$, which measures the error of the multicontinuum homogenization solution using the fine-grid solution as the reference. This quantifies the performance of the original multicontinuum homogenization method.
  • Type 2: $e_{2}^{(i)}(U_{h},U_{\textnormal{HMH}})$, which measures the error of the hierarchical multicontinuum homogenization solution relative to the fine-grid reference. This assesses the accuracy of the hierarchical approach.
  • Type 3: $e_{2}^{(i)}(U_{\textnormal{MH}},U_{\textnormal{HMH}})$, which quantifies the difference between the hierarchical solution and the standard multicontinuum homogenization solution. This isolates the error introduced by the hierarchical algorithm itself.
@@ 4.1 Example 1
We begin with a layered medium characterized by a high-contrast function $k$. The logarithm of $k$ together with the reference solution is displayed in Figure 5, while the corresponding three-level hierarchical structure ($L=3$) is illustrated in Figure 4. CPU timings for MH and HMH at one macropoint and its neighbors are reported in Table 4 for $H=1/12$ with oversampling of $l=5$ layers. The hierarchical approach achieves a reduction of nearly half in cost (101.21 seconds for MH vs. 46.9 seconds for HMH). Table 2 reports three error measures. The first two, corresponding to the two upscaling methods, decrease consistently as the coarse mesh is refined. The third, which quantifies the difference between MH and HMH, does not exhibit convergence but stays at a negligible level for all tested mesh sizes. The hierarchical method delivers accuracy comparable to MH, with differences remaining negligible; for instance, at $H=1/24$, both methods produce errors on the order of $10^{-2}$, differing by only $10^{-7}$. The average solutions of the reference, MH, and hierarchical approaches are plotted in Figure 6, confirming both accuracy and robustness.
Macropoints DOF CPU (s)
MH 16 $\mathcal{O}\left((\frac{1}{h})^{2}\right)$ 101.21
HMH 1 $\mathcal{O}\left((\frac{1}{h})^{2}\right)$ 46.90
3 $\mathcal{O}\left((\frac{2}{h})^{2}\right)$
12 $\mathcal{O}\left((\frac{4}{h})^{2}\right)$
Table 1: Comparison of DOF and CPU time of cell problems for MH and HMH methods in Example 1.
@@ 12 $\mathcal{O}\left((\frac{4}{h})^{2}\right)$
1/12 5 8.34e-02 8.04e-02 8.34e-02 8.04e-02 1.90e-06 2.12e-06
1/24 7 1.92e-02 1.74e-02 1.92e-02 1.74e-02 1.02e-07 4.22e-08
1/48 8 2.39e-03 2.32e-03 2.39e-03 2.32e-03 1.16e-07 1.16e-07
@@
Figure 5: High-contrast part $\log(k)$ (left) and reference solution (right) in Example 1.
(a) $U_{h}^{(1)}$
(b) $U_{h}^{(2)}$
(c) $U_{\textnormal{MH}}^{(1)}$
(d) $U_{\textnormal{MH}}^{(2)}$
(e) $U_{\textnormal{HMH}}^{(1)}$
(f) $U_{\textnormal{HMH}}^{(2)}$
Figure 6: Average solution in Example 1.
@@ 4.2 Example 2
The second test involves a crossed geometry for the coefficient $k$. As in Example 1, a three-level hierarchical approach ($L=3$) is used. The logarithm of $k$ and the reference solution appear in Figure 7, while Table 3 presents the errors for several coarse mesh sizes. The hierarchical method again performs on par with MH, with errors decreasing systematically under refinement. The discrepancy between the two approaches is insignificant compared to the homogenization error; for instance, the homogenization error can be as large as $10^{-3}$, whereas the difference between MH and HMH is only about $10^{-7}$ for $H=1/48$. Selected average solutions are shown in Figure 8, illustrating the consistency of the results.
Figure 7: High-contrast part $\log(k)$ (left) and reference solution (right) in Example 2.
@@
1/12 5 4.53e-02 4.53e-02 4.53e-02 4.53e-02 6.15e-06 6.35e-06
1/24 7 6.37e-03 6.23e-03 6.37e-03 6.23e-03 4.10e-07 3.77e-07
1/48 8 1.47e-03 1.41e-03 1.47e-03 1.41e-03 1.59e-07 1.55e-07
@@
(a) $U_{h}^{(1)}$
(b) $U_{h}^{(2)}$
(c) $U_{\textnormal{MH}}^{(1)}$
(d) $U_{\textnormal{MH}}^{(2)}$
(e) $U_{\textnormal{HMH}}^{(1)}$
(f) $U_{\textnormal{HMH}}^{(2)}$
Figure 8: Average solution in Example 2.
@@ 4.3 Example 3
We next turn to a non-periodic high-contrast function $k$. Figure 9 depicts the logarithm of $k$ alongside the fine-grid solution. A two-level hierarchical approach ($L=2$) is adopted in this case. For $H=1/12$ with oversampling of $l=5$ layers, CPU timings at one macropoint and its neighbors are given in Table 4. The hierarchical strategy reduces the computational cost by nearly a factor of two (26.42 seconds for MH vs. 12.01 seconds for HMH). Errors for two coarse mesh sizes are listed in Table 5. While the homogenization error decreases with mesh refinement, the hierarchical error does not exhibit the same convergence; nevertheless, it remains within a reasonable range, at roughly $10\%$ for the tested meshes. Figure 10 shows the average solutions from the reference, MH, and hierarchical methods, which align well with the numerical observations.
Macropoints DOF CPU (s)
MH 4 $\mathcal{O}\left((\frac{1}{h})^{2}\right)$ 26.42
HMH 1 $\mathcal{O}\left((\frac{1}{h})^{2}\right)$ 12.01
3 $\mathcal{O}\left((\frac{2}{h})^{2}\right)$
Table 4: Comparison of DOF and CPU time of cell problems for MH and HMH methods in Example 3.
@@ 3 $\mathcal{O}\left((\frac{2}{h})^{2}\right)$
1/12 5 4.71e-02 4.69e-02 1.10e-01 9.68e-02 6.72e-02 5.44e-02
1/24 7 1.28e-02 1.06e-02 7.46e-02 6.24e-02 7.73e-02 6.24e-02
@@
Figure 9: High-contrast part $\log(k)$ (left) and reference solution (right) in Example 3.
(a) $U_{h}^{(1)}$
(b) $U_{h}^{(2)}$
(c) $U_{\textnormal{MH}}^{(1)}$
(d) $U_{\textnormal{MH}}^{(2)}$
(e) $U_{\textnormal{HMH}}^{(1)}$
(f) $U_{\textnormal{HMH}}^{(2)}$
Figure 10: Average solution in Example 3.
@@ 4.4 Example 4
The fourth example considers another non-periodic configuration of $k$. A two-level hierarchical approach ($L=2$) is employed. Figure 11 illustrates the distribution of the two continua together with the reference solution. The errors for different coarse mesh sizes are summarized in Table 6. As the coarse mesh is refined, the homogenization error decreases, whereas the hierarchical error remains within an acceptable range (approximately $10\%$). Figure 13 displays the average solutions of the fine-grid, MH, and hierarchical approaches, showing consistent agreement among the three.
Figure 11: High-contrast part $\log(k)$ (left) and reference solution (right) in Example 4.
Figure 12: High-contrast part $k$ (left) and reference solution (right) in Example 5.
@@
1/12 5 6.26e-02 6.41e-02 1.20e-01 1.09e-01 6.23e-02 4.89e-02
1/24 7 1.92e-02 2.26e-02 8.49e-02 7.46e-02 7.66e-02 6.52e-02
@@
(a) $U_{h}^{(1)}$
(b) $U_{h}^{(2)}$
(c) $U_{\textnormal{MH}}^{(1)}$
(d) $U_{\textnormal{MH}}^{(2)}$
(e) $U_{\textnormal{HMH}}^{(1)}$
(f) $U_{\textnormal{HMH}}^{(2)}$
Figure 13: Average solution in Example 4.
4.5 Example 5
Finally, we test a three-dimensional case. The high-contrast function $k$ and the fine-grid solution are presented in Figure 12. CPU timings for MH and HMH are listed in Table 7, demonstrating that the hierarchical strategy reduces the cost by more than half (460.22 seconds for MH vs. 211.18 seconds for HMH). In this example we report results only for $H=1/10$, with the errors summarized in Table 8. The accuracy is satisfactory, with an error of about $7\%$, and the hierarchical approach achieves nearly machine precision when compared with MH. These findings confirm that the proposed method provides a reliable and efficient approximation also in three dimensions.
Macropoints DOF CPU (s)
MH 8 $\mathcal{O}\left((\frac{1}{h})^{3}\right)$ 460.22
HMH 1 $\mathcal{O}\left((\frac{1}{h})^{3}\right)$ 211.18
7 $\mathcal{O}\left((\frac{2}{h})^{3}\right)$
Table 7: Comparison of DOF and CPU time of cell problems for MH and HMH methods in Example 5.
$H$ $l$ Type 1 Type 2 Type 3
$e_{2}^{(1)}$ $e_{2}^{(2)}$ $e_{2}^{(1)}$ $e_{2}^{(2)}$ $e_{2}^{(1)}$ $e_{2}^{(2)}$
1/10 4 7.21e-02 7.82e-02 7.21e-02 7.82e-02 1.17e-16 8.43e-16
Table 8: Error comparison by different coarse mesh in Example 5.
(a) $U_{h}^{(1)}$
(b) $U_{h}^{(2)}$
(c) $U_{\textnormal{MH}}^{(1)}$
(d) $U_{\textnormal{MH}}^{(2)}$
(e) $U_{\textnormal{HMH}}^{(1)}$
(f) $U_{\textnormal{HMH}}^{(2)}$
Figure 14: Average solution in Example 5.
@@ Acknowledgement
Wei Xie acknowledges support from the China Scholarship Council (CSC, Project ID: 202308430231) for funding during the visit to Nanyang Technological University (NTU). Sincere thanks to NTU for providing an outstanding working and learning environment that greatly advanced the research and professional development. Viet Ha Hoang is supported by the Tier 2 grant T2EP20123-0047 awarded by the Singapore Ministry of Education. Yin Yang is supported by the grant Scientific Research In-novation Capability Support Project for Young Facult (No. SRICSPYF-BS2025132), National Natural Science Foundation of China Project (No. 12571469), the Project of Scientific Research Fund of the Hunan Provincial Science and Technology Department (No. 2024JJ1008), the 111 Project (No. D23017), and Program for Science and Technology Innovative Research Team in Higher Educational Institutions of Hunan Province of China.